### v3

@@ 3.2.3 Evaluation on Code Generation Task
3.2.4 Evaluation on Other Code-related Task
We score WaveCoder with state-of-the-art Code LLMs on HumanEvalPack in Table 4, 5 and use the result reported in Octopack [21] when avaliable. Table 4 presents the results of WaveCoder on code repair task and Table 5 presents the results of WaveCoder on Code summarization task, highlighting the the following salient observations:
WaveCoder outperforms all open source models on other code-related task with similar degrees of fine-tuning. Using Starcoder as the foundation model, our WaveCoder-SC outperforms both WizardCoder and OctoCoder across all programming languages on HumanEvalFix and HumanEvalExplain benchmarks. Even when compared with other foundation models, our model consistently demonstrates superior performance over all previously fine-tuned models.
Enhancing the refinement and diversity of data can significantly improve the effectiveness of instruction tuning in multi-task scenarios. By refining the data and classifying the instructions to 4 different code-related downstream tasks, our models achieve a surprising performance level on both evaluation scenarios. Notably, WaveCoder-DS, utilizing only 6.7B parameters, approximates the performance of GPT-4 on HumanEvalFix quite closely.
@@ 3.2.4 Evaluation on Other Code-related Task
3.3 Comparison with CodeAlpaca
CodeAlpaca dataset contains 20K multi-task instruction-following data generated by the techniques in the self-instruct [12]. To ensure a fair and multidimensional comparison, we randomly sampled 1K and 5K from both datasets (CodeAlpaca and CodeOcean), set the same set of training hyper-parameters set (epoch = 3, learning rate = 1e-4, LoRA rank = 8) and used the same training prompts. To prevent overfitting, we use Low-Rank Adaption (LoRA) [35] for fine-tuning if the size of instruction-follow training dataset is less than 5K and perform full fine-tuning on whole 20K dataset. Table 4, 5 present the result of CodeOcean on HumanEval and other code-related benchmarks highlighting the following salient observations:
1) After being fine-tuned with 1K, 5K and 20K of instructional data respectively, the performance of base model improves significantly on HumanEval shown in Figure 1. Taking Starcoder as the base model, CodeOcean surpasses the CodeAlpaca (44.9% vs 41.7%, 45.7% vs 48.1% and 47.0% vs 50.5%) shown in Figure 1 (a), which emphasizes the effectiveness of our LLM-based Generator-Discriminator Framework on refining instruction data. As shown in Figure 1 (b), The results of different base models on CodeOcean surpasses the results on CodeAlpaca, which emphasizes the effectiveness of CodeOcean dataset in enhancing the instruction-following ability of the base model.
4 Ablation and Analysis
To evaluate the the effectiveness of each code-related task appearing in training data and the relationship between different tasks, we conduct an ablation study about the task type of instruction data. Using DeepseekCoder-Base-6.7B as our base model, we have the following observations from Table 6:
Refined instruction data can significantly improve the generalization ability of pre-trained models without a tradeoff. As shown in Table 6, incorporating all 4 code-related tasks into training data, WaveCoder-DS-6.7B achieves the best performance on benchmark of all tasks. For example, the participation of the Code Repair task yields a considerable average improvement of 33.7% absolute for HumanEvalFix without any significant decline in other tasks, and even improved by 3.1% absolute for HumanEval benchmark.
Different tasks can promote each other so that the model can show a generalization ability. From Table 6, we can observe that any combination of three tasks resulted in a lower score than all tasks. For example, the addition of the code summarization task offers a modest yet significant average improvement on all benchmarks. Moreover, the absence of any task will cause the score of HumanEval to drop, which also reflects the mutual promotion between different tasks.
@@ 4 Ablation and Analysis
-Without Code Repair | ✔ | ✔ | ✔ | ✘ | 60.9 (-3.1) | 15.7 (-33.7) | 41.2 (-0.1) |
-Without Code Generation | ✘ | ✔ | ✔ | ✔ | 53.6 (-10.4) | 47.4 (-2.0) | 40.5 (-0.8) |
-Without Code Translation | ✔ | ✔ | ✘ | ✔ | 60.9 (-3.1) | 49.3 (-0.1) | 41.6 (+0.3) |
-Without Code Summarization | ✔ | ✘ | ✔ | ✔ | 61.5 (-2.5) | 45.6 (-3.8) | 28.4 (-12.9) |
@@ 5 Related Work
5.1 Code Large Language Models
Numerous researchers have proposed their Code Large Language Models (LLMs), including CodeGen [36], CodeT5[37], StarCoder [6], CodeLLaMa [7], and Deepseek-Coder [22] within coding domain. By leveraging a comprehensive pre-training process with a vast amount of code knowledge from scratch, these Code LLMs often exceed the performance of super LLMs in many scenarios. In a bid to generate output more efficiently and assist in resolving programming issues, instruction-based models like InstructCodeT5+ [38] and WizardCoder [14] have been introduced. When drafting this paper, we noticed that our contemporaneous work Magicoder [34] has been released, so we conduct a brief comparative analysis in Section 3.
5.2 Instruction Tuning
In order to unleash the potential of the foundation LLMs, instruction tuning is proposed to fine-tune the LLM with multi-task datasets which is presented in task descriptions by natural language. For example, OpenAI fine-tuned its GPT-3 [3] model with a diverse instruction dataset written by human annotators, the instruction-tuned model can effectively align the intention of human users. This work is the indispensable part of the InstructGPT [17], which is the foundation of the well-known ChatGPT. Flan-PaLM 540B [10] is instruction-tuned on 1.8K tasks and finds that the diversity of instruction dataset can significantly improve the performance of the instruction-tuned language model. Alpaca [12] adopts the self-instruct method, which leverages the few shot ability of ChatGPT to generate data for training. For Code LLMs, InstructCodeT5+ [38], WizardCoder [14], Pangu-coder2 [32], are recent recent models trained with coding instructions. However, all the instruction data they used is from CodeAlpaca, which is not refined enough.
@@ 6 Conclusion
This paper studies the effect of multi-task instruction data on enhancing the generalization ability of Code Large Language Model. We introduce CodeOcean, a dataset with 20K instruction instances on four universal code-related tasks. This method and dataset enable WaveCoder, which significantly imporove the generalization ability of foundation model on diverse downstream tasks. Our WaveCoder has shown the best generalization ability among other open-source models in code repair and code summarization tasks, and can maintain high efficiency on previous code generation benchmarks. In order to further improve the mono-task performance and generalization ability of the model, future work should focus on the interplay among different tasks and larger dataset.

### v4

@@ 2.2.2 LLM-based Generator-Discriminator Framework
| WaveCoder-Pro-6.7B |
59.1 | 56.7 | 54.2 | 45.1 | 45.7 | 34.1 | 49.2 |
| Deepseek-instruct-6.7B | 56.1 | 58.5 | 57.3 | 49.4 | 45.1 | 36.6 | 50.5 |
| Magicoder-S-DS | 56.1 | 55.4 | 58.5 | 51.2 | 45.7 | 35.3 | 50.3 |
| WaveCoder-Ultra-6.7B |
58.5 | 57.3 | 61.0 | 53.0 | 50.0 | 37.2 | 52.8 |
3 Experiments
3.1 Setup
Unlike the previous work Luo et al. (2024); Shen et al. (2023); Gunasekar et al. (2023) that mainly focus on code generation task, we generate about 20K dataset covers 4 common code-related tasks to enhance the geralization abilties of Code LLM. To obtain WaveCoder models, We choose StarCoder-15B, CodeLLaMa (7B and 13B), DeepseekCoder-6.7B as the base model and fine-tune all the base model for 3 epochs using NVIDIA A100-80GB GPU. For StarCoder-15B, CodeLLaMa-7B and CodeLLaMa-13B, we set the global batch size to 256 using Tensor Parallel and set the initial learning rate at 2e-5. For DeepseekCoder-6.7B, we set the global batch size to 512 using the Fully Sharded Data Parallel (FSDP) module from Pytorch and set the initial learning rate at 5e-5.
Benchmarks and Baselines. To ensure a thorough assessment of the model’s generalization ability, we score our model on three code benchmarks across different code related tasks: HumanEval Chen et al. (2021), MBPP Austin et al. (2021) and HumanEvalPack Muennighoff et al. (2024), as illustrated in Appendix D.
Proprietary Models. We present the self-reported results from an array of SoTA LLMs, including ChatGPT (gpt-3.5-turbo), GPT-4. If not reported, we use the results from Octopack Muennighoff et al. (2024) or evaluate by ourselves.
Open Source Models. To ensure an equitable comparison, we opted to select models that have been trained with the similar amount of instruction instances for our comparative analysis.
SoTA Open Source Models. We compared WaveCoder-6.7B with the SoTA open source Code LLM, includes Magicoder-S-DS Wei et al. (2023) and DeepseekCoder-instruct-6.7B Wei et al. (2023) on a wide range of code-related tasks. All the result of SoTA open source models is presented from EvalPlus. Liu et al. (2023) If not reported, we evaluate it by ourselves.
@@ 3.1 Setup
| WaveCoder-Pro-6.7B |
53.0 | 43.3 | 54.9 | 34.1 | 42.7 | 20.0 | 41.3 |
| Magicoder-S-DS | 60.3 | 46.3 | 54.3 | 38.4 | 48.1 | 29.2 | 46.1 |
| Deepseek-instruct-6.7B | 62.2 | 54.3 | 61.0 | 39.6 | 55.5 | 33.5 | 51.0 |
| WaveCoder-Ultra-6.7B |
56.7 | 50.0 | 54.3 | 34.8 | 51.2 | 36.6 | 47.3 |
3.2 Result
Evaluation on Code Generation Task. HumanEval and MBPP are two representative benchmarks for code generation task, as illustrated in Appendix D. Table 3 shows the pass@1 score of different LLMs on both benchmarks. From the results, We have the following observations:
1) WaveCoder-Pro-6.7B outperforms other open source models with only 6.7B parameters and 20K instruction data. Trained with GPT-4 enhanced CodeSeaXDataset dataset, WaveCoder-Pro-6.7B achieve 72.0% pass@1 and on HumanEval and 63.6% on MBPP, surpassing all open source models but still behind proprietary models and the SoTA open source models.
2) Refined and diverse instruction data can significantly improve the efficiency of instruction tuning. As delineated in Table 3, WaveCoder demonstrates commendable performance, utilizing a dataset comprising merely about 20K Instruction Tuning Data (InsT Data), which positions it on an equal footing with its contemporaries. Despite a discernible shortfall in the code generation benchmarks relative to WizardCoder (50.5 vs 57.3) and Magicoder (64.0 vs 66.5), it is imperative to consider the substantial disparity in the volume of training data. Moreover, it is observed that WaveCoder-pro-6.7B significantly outperforms Magicoder-DS-6.7B (72.0 vs 66.5), demonstrating the effectiveness of data quality and diversity in instruction tuning.
Evaluation on Other Code-related Task. We score WaveCoder with state-of-the-art Code LLMs on HumanEvalPack Muennighoff et al. (2024) in Table 4 and Table 5, highlighting the the following salient observations:
1) WaveCoder models outperform all open source models on other code-related task. Building upon Starcoder, our proposed WaveCoder-SC has exhibited exceptional performance, transcending the capabilities of both WizardCoder and OctoCoder as evidenced by the HumanEvalFix (33.0 vs 25.7 vs 27.0) and HumanEvalExplain (30.8 vs 27.5 vs 24.5) benchmarks, which is also shown in other base models. Notably, WaveCoder-DS-6.7B achieves 49.4% average pass@1 score on HumanEvalFix and 41.3% on HumanEvalExplain, surpassing all open source models and demonstrating strong generalization capabilities in multi-task scenarios.
2) The enhancement in data refinement and diversification can markedly bolster the efficacy of instruction tuning in multi-task scenarios. Such data refinement, coupled with the categorization of instructions into four code-related tasks, has propelled our models to reach an unforeseen generalization capabilities in various code-related tasks. Remarkably, our WaveCoder-DS-6.7B model outperforms GPT-4 (49.4 vs 47.8) on HumanEvalFix, thereby underscoring the potential of smaller models to achieve near-parity with parameter-heavy models when optimized efficiently.
WaveCoder-Ultra-6.7B. Inspired by Magicoder-S-DS-6.7B Wei et al. (2023), we combine CodeSeaXDataset with WaveCoder-evol-codealpaca to a 130K dataset. Fine-tuned with this combination of two datasets, we obtain WaveCoder-Ultra-6.7B. As illustrated in Table 3, 4, 5, WaveCoder-Ultra-6.7B has the state-of-the-art generalization abilities on a wide range of code-related tasks, which highlights the significance of our CodeSeaXDataset dataset again and demonstrates the potential of larger datasets.
@@ 3.2 Result
| -Without Repair | ✔ | ✔ | ✔ | ✘ | 60.9 (-3.1) | 15.7 (-33.7) | 41.2 (-0.1) |
| -Without Generation | ✘ | ✔ | ✔ | ✔ | 53.6 (-10.4) | 47.4 (-2.0) | 40.5 (-0.8) |
| -Without Translation | ✔ | ✔ | ✘ | ✔ | 60.9 (-3.1) | 49.3 (-0.1) | 41.6 (+0.3) |
| -Without Summarization | ✔ | ✘ | ✔ | ✔ | 61.5 (-2.5) | 45.6 (-3.8) | 28.4 (-12.9) |
4 Ablation and Analysis
4.1 Ablation of Code-related Tasks
To explore the relationship between different tasks, we conduct an ablation study about the task type of instruction data. Using DeepseekCoder-Base-6.7B as our base model and initial 20K CodeSeaXDataset data as our base dataset, we have the following observations from Table 6:
1) Refined instruction data can significantly improve the generalization ability of pre-trained models without a tradeoff. As shown in Table 6, incorporating all 4 code-related tasks into training data, WaveCoder-DS-6.7B achieves the best performance on benchmark of all tasks. For example, the participation of the Code Repair task yields a considerable average improvement of 33.7% absolute for HumanEvalFix without any significant decline in other tasks, and even improved by 3.1% absolute for HumanEval benchmark.
2) Different tasks can promote each other so that the model can show a generalization ability. From Table 6, we can observe that any combination of three tasks resulted in a lower score than all tasks. For example, the addition of the code summarization task offers a modest yet significant average improvement on all benchmarks. Moreover, the absence of any task will cause the score of HumanEval to drop, which also reflects the mutual promotion between different tasks.
4.2 Discussion about Data Leakage
In this section, we explore the potential leakage through three instruction datasets about code (i.e. Code Alpaca, CodeSeaXDataset, Magicoder-evol-codealpaca). To ensure an accurate analysis, we employ SoTA embedding model GTE-Large Li et al. (2023b) to encode the canonical code in test benchmarks and all code in training set. Subsequently, we find the nearest neighbour in train set for each questions in test benchmark. As illustrated in Figure 3, CodeSeaXDataset has the lower average cosine similarity than other datasets. Figure 6 in Appendix presents two examples about the data leakage in these training set. Moreover, we analyze all benchmarks and notice a serious data leakage issue between HumanEval and Magicoder-evol-codealpaca dataset. Therefore, we decontaminate Magicoder-evol-codealpaca for each evaluation problem in HumanEval and obtain WaveCoder-evol-codealpaca. As illustrated in Figure 3, WaveCoder-evol-codealpaca has lower similarity than Magicoder-evol-codealpaca.
@@ 5 Related Work
Instruction Tuning. Recent studies, such as FLAN Wei et al. (2022), ExT5 Aribandi et al. (2022), and FLANT5 Chung et al. (2022), have underscored the efficacy of integrating diverse tasks within training process to bolster the adaptability of pre-trained models for downstream tasks. Specifically, Flan-PaLM 540B’s Chung et al. (2022) instruction-tuning over 1.8K tasks has demonstrated that a widespread and versatile enhanced instruction dataset markedly enhances language model performance. InstructGPT Ouyang et al. (2022), with its incorporation of premium instruction data crafted by human annotators, has shown significant promise in aligning model outputs with user intents, prompting further investigation into instruction-tuning mechanisms. Additionally, Stanford Alpaca Taori et al. (2023) has innovatively employed GPT-generated instruction data via self-instruct Wang et al. (2023a) for instruction tuning process. WizardLM Xu et al. (2024) has built upon these advancements by applying the evol-instruct methodology, collectively illuminating the transformative impact of instruction tuning on the overall capabilities of LLM.
Code Large Language Models. Recent advancements in code generation have been propelled by Code LLMs such as CodeGen Nijkamp et al. (2022), CodeT5 Wang et al. (2021), StarCoder Li et al. (2023a), CodeLLaMa Roziere et al. (2023) and Deepseek-Coder Guo et al. (2024), which benefit from extensive pre-training on expansive code corpora. Efforts to further enhance efficiency and problem-solving capabilities have led to the development of instruction-tuned models like InstructCodeT5+ Wang et al. (2023b), WizardCoder Luo et al. (2024), Pangu-coder2 Shen et al. (2023), However, all the instruction data they used is from Code Alpaca, which is not refined enough in the context of multi-task environment, which drives us to propose new methods for instruction data generation. Concurrently, with the release of our contemporaneous work Magicoder Wei et al. (2023), we offer a concise analysis in Section 3.
@@ 6 Conclusion
This paper presents WaveCoder, a Code LLM fine-tuned with widespread and versatile enhanced instruction data. By enabling language models to effectively tackle complex code-related tasks, our approach demonstrates the potential of integrating multiple code-related tasks into instruction tuning for Code LLM and generating high-quality and diverse instruction data for specific task requirements in multi-task scenarios. WaveCoder achieves state-of-the-art generalization performance on different code-related tasks surpassing existing open source Code LLMs. Furthermore, our analysis of the relationship of different tasks provides valuable insights for future research, paving the way for more extensive code-related tasks and larger dataset.
Limitations
We present WaveCoder and propose a data generation method which can stably generate high-quality and diversity instruction data from open source dataset in multi-task scenario. One limitation of our work is that the training dataset we used only includes 19,915 instructions, which produces limited enhancements to the model. As illustrated Section 3, we expand the training dataset to a larger amount and the resulted model still have significant improvement. Therefore, future work should focus on more code-related task types and larger dataset.
Ethics Statement
We constructed our CodeSeaXDataset dataset from open source code. For each code snippet we used, we are committed to adhering to the terms of its license, which includes proper attribution ensuring that any modifications or derivative works are also shared under the compatible terms. Moreover, we notice a serious data leakage issue in the Magicoder-evol-codealpaca dataset. To ensure a fair comparison, we remove three nearest neighbours of each question in test benchmark from train set. However, if all similar samples are accidentally removed, the integrity of the data will be damaged, which is harmful to model training. Therefore, this phenomenon should be attributed to the fact that the problems in the current test benchmarks are some of the most basic algorithm logic. To this end, we call for more comprehensive and complex test benchmarks for Code LLMs which will not easily cause data leakage problem.
Appendix A Prompt
Followed Alpaca Taori et al. (2023), we set the fine-tuning prompt as follows:
Appendix B Examples of the LLM-based Generator-Discriminator framework
| LLM-based Generator-Discriminator framework |
| Input: |
| The Python code calculates the area of a circle given its radius. |
| Output: |
| task_name: Calculate Circle Area |
| instruction: Write a Python function that calculates the area of a circle given its radius. |
| information: The formula to calculate the area of a circle is A = pi * , where A is the area and r is the radius. |
| solution: |
| import math def area_of_circle(radius): return math.pi * radius ** 2 |
| Analysis: - Step 1: Check the Instruction: The programming language should be specified in the instruction. <answer: yes, The instruction mentions "Write a Python function," which clearly specifies the programming language as Python.> |
| - Step 2: Check the Solution: |
| 1. Solution is related to instruction and information. Solution is the specific resolution to instruction and information. <answer: yes, The solution provided calculates the area of a circle given its radius, which is directly related to the instruction and the information provided.> |
| 2. In the solution, it should only contain the code and comments within the code. There should be no explanations provided outside the code. <answer: yes, The solution only contains the code without any explanations outside the code. Comments within the code are not present, but the code is straightforward and easy to understand.> |
| 3. The code that contains algorithmic logic in solution should not often be buried inside complex or poorly documented functions, making them difficult to follow or learn from. <answer: yes, The code that contains algorithmic logic in the solution should not often be buried inside complex or poorly documented functions, making them difficult to follow or learn from.> |
| 4. The code should import the required necessary libraries. <answer: yes, The code imports the necessary math library, which is required for calculating the area of a circle using the mathematical constant pi.> |
| - Overall answer: yes |
| - Reasons: All the requirements are met as per the given rules. The programming language is specified, and the solution is related to the instruction and information. The code is simple, easy to follow, and imports the necessary library. The solution provided is a good example of a well-structured and concise code that meets the task’s specifications. |
Appendix C Comparison with CodeAlpaca
CodeAlpaca dataset contains 20K multi-task instruction-following data generated by the techniques in the self-instruct Taori et al. (2023). To ensure a fair and multidimensional comparison, we randomly sampled 1K and 5K from both datasets (CodeAlpaca and CodeSeaXDataset), set the same set of training hyper-parameters set (epoch = 3, learning rate = 1e-4, LoRA rank = 8) and used the same training prompts. To prevent overfitting, we use Low-Rank Adaption (LoRA) Hu et al. (2022) for fine-tuning if the size of instruction-follow training dataset is less than 5K and perform full fine-tuning on whole 20K dataset.
1) After being fine-tuned with 1K, 5K and 20K of instructional data respectively, the performance of base model improves significantly on HumanEval shown in Figure 5. Taking Starcoder as the base model, CodeSeaXDataset surpasses the CodeAlpaca (44.9% vs 41.7%, 45.7% vs 48.1% and 47.0% vs 50.5%) shown in Figure 5 (a), which emphasizes the effectiveness of our method on refining instruction data. As shown in Figure 5 (b), The results of different base models on CodeSeaXDataset surpasses the results on CodeAlpaca, which emphasizes the effectiveness of CodeSeaXDataset dataset in enhancing the instruction-following ability of the base model.
Appendix D Evaluation Benchmark